OpenStudy (anonymous):

Need help with this one!!! A panel of judges must consist of four women and three men. A list of potential judges includes six women and five men. How many different panels could be created from this list?

OpenStudy (anonymous):

(6 women + 5 men) = 11 ppl to choose from, (4 women + 3 men) = 7 ppl chosen This is a combination, order doesn't matter. Otherwise it would be a permutation (sp?). I think... Men : 5 nCr 3 = 10 Women : 6 nCr 4 = 15 Sum the two for 25 combinations total? Can somebody please check me on this? I don't really feel confident this is necessarily the answer, sorry. :-/

OpenStudy (anonymous):

the posible answers for this question are a) 30 b) 150 c)216 so 25 is not the answer :/ but thanks for trying!

OpenStudy (anonymous):

Thanks @palomamtz15 all I recalled was this forumula: \[\frac{n!}{(n-r)!}\] where n is the number of things to choose from, and you choose r of them and this fact: If the order doesn't matter, it is a Combination. If the order does matter it is a Permutation.

OpenStudy (anonymous):

10*15 = 150 = answer?

OpenStudy (anonymous):

o_O

OpenStudy (anonymous):

THAT was the ANSWER!! Thanks a lot! :D

OpenStudy (anonymous):

Ah I see... That's the permutation formula above, the combination formula is: \[ nCr = \frac{n!}{r!(n-r)!} \] \[ nCr = \frac{5!}{3!(5-3)!} = 10\]

OpenStudy (anonymous):

And there's the math that is going on in my calculator :P

OpenStudy (anonymous):

:) thanks for the help! I really need that answer!

OpenStudy (anonymous):

5! = 1*2*3*4*5 = 120

OpenStudy (anonymous):

needed*